# Towards Assessing Medical Ethics from Knowledge to Practice

### v1

@@ 3.1 Data Curation and Annotation
Note. In (a) Data Curation phase, entities highlighted in blue represent GPT-4o. “SOTA LLMs” refers to GPT-4.1, Gemini 2.5 Pro, and Claude 4 Sonnet.
Figure 1: Benchmarking workflow of PrinciplismQA.
3.1.1 Knowledge MCQA Curation
The knowledge MCQAs were curated using a systematic and rigorous multi-stage process. Initially, we introduced GPT-4o to segment the content from 350 authoritative textbooks, extracting 93k Contents of Interest (COIs) identified as exam-worthy from tables of contents. Subsequently, GPT-4o acted as an LLM-as-an-exam-creator, curating four-option MCQAs directly from these COIs, intentionally generating plausible distractors from the original context wherever possible. This procedure yielded approximately 90k curated questions.
A preliminary MCQA quality assessment utilized 3 state-of-the-art (SOTA) baseline models, GPT-4.1, Gemini 2.5 Pro, and Claude 4 Sonnet, filtering out overly simplistic questions to retain approximately 5,000 high-quality questions. After further excluding questions that were contextually ambiguous or misaligned with international consensus standards, we retained about 2,500 candidate questions. GPT-4o was again employed for pseudo-labeling ethical principles associated with each MCQA in a one-hot encoding format. The criteria of Principlism labeling is listed in Table 1 introduced in Section 2.1.
3.1.2 Open-Ended Practice Question Curation
We sourced a total of 1,466 medical ethics case analysis articles from the “CASE AND COMMENTARY” section of the AMA Journal of Ethics website. Articles in this section are typically authored by one or more medical, legal, or ethics experts, beginning with a concrete clinical or ethical scenario and followed by multi-perspective ethical analysis and practical recommendations. From these, we selected 702 articles exhibiting a clear “case–commentary–discussion” structure for the subsequent extraction of open-ended questions.
Leveraging LLMs, we systematically extracted author-raised questions as well as passages articulating “considerations” and “decision-making recommendations” from peer-reviewed expert commentaries relevant to each case. These passages were further decomposed into multiple key-point sentences, each representing an essential element of ethical reasoning that reflects consensus human values. These keypoints collectively formed the checklist-style reference answers for the open-ended questions. Furthermore, each keypoint was annotated to indicate which of the six ACGME Core Competencies it assessed, enabling precise mapping between answer content and competency domains. To address copyright requirements, the original case descriptions were moderately rephrased based on the extracted questions.
Each step of this pipeline was reviewed through multi-round discussion among three SOTA LLMs, with consensus reached before proceeding. If consensus could not be achieved among the three LLMs within three rounds of discussion, the corresponding question was excluded from the dataset. Figure 1 illustrates the detailed workflow for generating open-ended practice questions and corresponding reference answers.
3.1.3 Annotation
A panel of twelve medical experts conducted human-in-the-loop verification and refinement over a period of three months. This process involves three levels: (1) question quality, (2) reference answer soundness, and (3) correctness of Principlism classification for both questions and answers. The panel comprised four practicing physicians and eight medical postgraduates, all of whom had completed formal coursework in Medical Ethics within the past six months. These experts, drawn from diverse medical departments, were evenly distributed by gender.
During the expert review stage, questions that were erroneous or of low quality were removed. Additionally, we ensured that the reference answer checklists for open-ended questions objectively represented consensus human values in medical ethics, rather than reflecting only the original authors’ subjective opinions. After this rigorous manual review, we retained 87.3% of MCQAs and 96.4% of open-ended questions. Notably, among the retained open-ended questions, only 2.8% of the keypoints required expert revisions, underscoring the effectiveness of our curated question generation pipeline.
After this comprehensive process, we curated a high-quality final set of 2,182 MCQAs and 1,466 open-ended practice medical ethics questions with validated reference answers, derived from 677 unique clinical ethics cases. Notably, a single case may correspond to multiple questions, as each case often encompasses several key ethical issues or decision points.
Table 4: Keypoint Distribution by Core Competency
Core Competency # of Keypoints
Interpersonal & Communication 1,490 (22.7%)
Medical and Ethical Knowledge 304 (4.6%)
Patient Care 836 (12.8%)
Practice-based Learning 75 (1.1%)
Professionalism 2,870 (43.8%)
Systems-based Practice 980 (15.0%)
Total 6,555
In addition, each keypoint in the reference answers for open-ended questions was annotated with the specific ACGME Core Competency it assessed, covering the critical areas defined by the ACGME—including Patient Care, Medical Knowledge (specifically medical and ethical knowledge), Practice-Based Learning and Improvement, Interpersonal and Communication Skills, Professionalism, and Systems-Based Practice (Edgar et al., 2020). As shown in Table 4, a total of 6,555 keypoints were annotated, enabling fine-grained analysis of LLMs’ strengths and weaknesses across different ethical and clinical competency domains.
3.2 Evaluation Protocols
Open-Ended Response Stepwise Evaluation Sample Case: Dr. Avery, an emergency medicine resident, treats Scott Daley, who is semi-conscious and mildly intoxicated after a head injury. Daley refuses further testing and wants to leave, but his decisional capacity is unclear. Dr. Avery faces the ethical question of whether restraint is justified for medical safety. Question 1: Is it ethically justifiable to override the patient’s wishes and restrain him to ensure medical safety if decisional capacity cannot be fully assessed? LLM-as-a-Judge Stepwise Evaluation Keypoint Rubric Sample Model Response Fragment Keypoint Score Partial decisional capacity but impaired understanding. Discusses uncertainty of capacity, but lacks detail on partial capacity. 0.5 Temporary detention justified if evaluation incomplete. Justifies temporary detention for safety. 1.0 Patient should be informed if detained. Mentions restraint as last resort, but not the importance of informing patient. 0.5 Respect autonomy if risk is understood; otherwise restrain. Focuses on risk of harm, not on autonomy if patient understands risks. 0.5 Address reversible impairments (alcohol, pain, etc.). Does not discuss mitigating reversible factors. 0.0 Practice Score: $(0.5+1.0+0.5+0.5+0.0)/5=0.5$ LLM-as-a-Judge Justification: The model partially addresses key ethical considerations but omits some crucial elements such as detailed discussion of partial capacity, explicit patient communication, and identification of reversible factors.
Figure 2: Illustrative stepwise evaluation of an open-ended response: Model outputs are aligned with keypoint rubrics, scored, and justified using the LLM-as-a-Judge protocol.
To ensure clarity and comparability, each question is presented to the models in a prompt that briefly specifies their role as a medical ethics expert. The same batch of model testing should be under consistent experimental settings, which are detailedly described in Section 4.1.
For the Knowledge MCQAs, models are required to select a single best answer from four options and provide a concise rationale grounded in medical ethics principles. Responses must be formatted strictly containing the selected answer and the reason, which minimizes ambiguity and facilitates automated evaluation. Model performance is measured primarily by accuracy, calculated as the proportion of correct responses, with no answer aggregation or exposure to reference answers.
Using expert-validated keypoint checklists as reference answers, we evaluate LLM performance on open-ended Practice questions with a systematic LLM-as-a-Judge protocol. For each case, model-generated answers are systematically compared against these keypoints, each of which represents a discrete and essential element of ethical reasoning. The evaluation LLM (GPT-4o), acting as an automated expert judge, receives the original question, keypoint checklist, and candidate answer in a structured prompt. Each keypoint is then scored independently using a three-level rubric:
  • 1.0 for complete and accurate coverage, with no errors or repeats;
  • 0.5 for partially correct or incomplete content, with minor omissions or slight redundancy;
  • 0.0 for missing, incorrect, or excessively redundant content.
The LLM-as-a-Judge outputs structured evaluation results with per-keypoint numerical scores and concise justifications for each decision. This process ensures consistency and scalability in evaluating complex, open-ended responses. A stepwise evaluation sample is depicted in Figure 2.
As shown in Figure 2, each question’s practice score is determined by its averaged keypoint scores, yielding a normalized value between 0 and 1 for fair comparison across cases of varying complexity.
To enable Principlism-specific analysis, we categorize each test item by its primary ethical principle. For both Knowledge and Practice questions, performance metrics are calculated separately for each principle. This approach enables systematic comparison of model strengths and weaknesses across ethical domains.
4 Experimental Study
@@ 4.1 Experiment Settings
To comprehensively assess Principlism-based ethical reasoning in both general-purpose and domain-specialized language models, we include a broad set of recent medical LLMs alongside general models(Chen et al., 2024; Christophe et al., 2024; Sellergren et al., 2025; Liu et al., 2024). Medical models are fine-tuned for healthcare contexts, allowing us to evaluate whether domain adaptation improves ethical sensitivity and Principlism coverage in clinical scenarios. Besides, representative closed-source LLM families, including ChatGPT(OpenAI, 2025), Claude 4(Anthropic, 2025), Qwen3(Yang et al., 2025), and Gemini 2.5(Comanici et al., 2025), were involved to benchmark the performance of widely used proprietary systems. Commonly used baseline models, LlaMA3.1(Dubey et al., 2024), Qwen 2.5(Qwen et al., 2025), and Gemma(Team et al., 2025), are evaluated both as general LLMs and as the bases for their corresponding medical model variants, enabling direct comparison between general and medical domain LLMs. All tested LLMs are listed in Table 5.
Table 5: Evaluated Models and Inference Methods
Model API Provider or HuggingFace Checkpoint
@@
All evaluations were conducted using a fixed sampling temperature of 0.1, regardless of whether the model was accessed via API or hosted locally. For all open-source models, the maximum sequence length was set to 2048 tokens. Each question was tested with a single response per model, with no answer aggregation. Open-source model inference was performed on four NVIDIA H20 GPUs (140GB each), using the original precision as provided by official HuggingFace checkpoints. The prompt constraints and evaluation metrics to be obtained are detailed in Section 3.2.
4.2 Results and Analysis
Table 6: Overall Performance of All LLMs on PrinciplismQA.
Model Knowledge Practice Overall
OpenAI o3 74.4 80.7 77.5
Qwen-Plus 70.0 73.3 71.6
Gemini 2.5 Flash 70.2 72.4 + 71.3 +
OpenAI o3-mini 73.3 67.2 70.2
Claude Sonnet 4 Thinking 70.0 67.5 + 68.7 +
DeepSeek-R1 68.0 + 66.6 + 67.3 +
Gemma3-27B 65.5 40.1 52.8
Gemma3-4B 59.5 42.8 51.1
GPT-4.1 74.7 70.8 72.7
Gemini 2.5 Flash Non-thinking 70.4 + 69.5 69.9
Claude Sonnet 4 70.0 66.6 68.3
Llama-3.1-70B 69.7 55.6 62.6
DeepSeek-V3 66.5 62.5 64.5
Qwen2.5-72B 69.8 53.5 61.7
Qwen2.5-7B 66.4 49.4 57.9
Llama-3.1-8B 58.4 48.5 53.5
HuatuoGPT-o1-72B 70.1 61.6 65.9
HuatuoGPT-o1-70B 67.5 61.3 64.4
Med42-70B 67.4 61.2 64.3
MedGemma-27B 64.4 64.3 64.3
HuatuoGPT-o1-7B 66.5 55.2 60.8
HuatuoGPT-o1-8B 55.4 56.4 55.9
MedGemma-4B 59.4 52.9 56.1
Med42-8B 60.5 49.6 55.1
Notes. The bold data are the most significant performance in the same category, while the underlined data are the weakest performance. The red color highlights the highest performance and the blue refers to the weakest. “+” denotes stronger performance of reasoning and chat variants within a model family. $\uparrow$” indicates metric improvement of a medical model compared to its general-domain baseline model. All subsequent tables follow these annotation conventions.
Takeaway 1: Ethical issues exist for every LLMs.
Takeaway 2: LLMs know ethics but struggles with practice.
Large Reasoning Model vs. Large Language Model in Ethics Across all evaluated models, state-of-the-art closed-source and general reasoning models demonstrated the strongest performance in medical ethics tasks. For example, o3 achieved the highest overall score of 77.5, with 74.4% Knowledge accuracy and an 80.7 performance on Practice, while GPT-4.1 led among chat models with an overall score of 72.7—both outperforming cutting-edge Practice performance and specialized medical models. Furthermore, reasoning-focused variants, such as gemini-2.5-flash, claude-sonnet-4, and deepseek, consistently surpassed their chat-oriented counterparts in Practice scenarios. These results suggest that models with stronger foundational and reasoning capabilities are better equipped to handle complex, non-standardized ethical dilemmas in the medical domain.
Takeaway 3: Reasoning helps ethics.
4.3 Fine-grained Analysis
Takeaway 5: LLMs struggle most with beneficence; fine-tuning helps.
Table 7: Principlism-Specific Performance of All Models on PrinciplismQA.
Model Autonomy Nonmaleficence Beneficence Justice
Claude Sonnet 4 Thinking 0.699 0.681 0.690 0.777 0.675 0.726 0.464 0.682 0.573 0.800 0.659 0.730
Gemini 2.5 Flash Non-thinking 0.687 0.700 0.694 0.736 0.705 0.720 0.491 0.701 0.596 0.812 0.671 0.741
Note. “Know.” and “Prac.” are the abbreviations of “Knowledge” and “Practice”.
Note. “ICS.”, “MK.”, “PBLI.”, “PC.”, “Pf”, and “SBP.” are the abbreviations of “Interpersonal and Communication Skills”, “Medical and Ethical Knowledge”, “Practice-Based Learning and Improvement”, “Patient Care”, “Professionalism”, and “Systems-Based Practice”.
Figure 3: Competency-specific open-ended question performance comparisons: (a) by model category, (b) between medical LLMs and their baseline models.
4.4 Trustworthiness of LLM-as-a-Judge
We evaluated the reliability of the LLM-as-a-Judge protocol (using GPT-4o) for open-ended questions by comparing its keypoint-level scores with those of three clinical experts on 480 sampled question–response pairs (1,516 keypoints). The inter-rater reliability (ICC ) among human experts was 0.67, reflecting good grading consistency despite inherent subjectivity. The ICC between GPT-4o and the human mean was 0.71, indicating that LLM-as-a-Judge achieves grading consistency comparable to human experts (see Table 8).
These results show that LLM-as-a-Judge can reliably substitute for human grading in open-ended medical ethics assessments.
Table 8: Inter-rater reliability (ICC) for LLM-as-a-Judge
Grader Comparison ICC
Human–Human (3 experts) 0.67
LLM-as-a-Judge (GPT-4o) vs. Human Mean 0.71
5 Conclusion
We introduce PrinciplismQA, a comprehensive benchmark designed to systematically evaluate the ethical reasoning of Large Language Models in medicine. Grounded in the core principles of medical ethics, it uniquely combines knowledge-based questions with practice-oriented case studies.
Our extensive evaluation reveals several critical insights. First, a significant “knowledge-practice gap” exists across all models, which especially struggle to dynamically apply ethical principles to real-world scenarios. Most LLMs perform poorly in dilemmas concerning “beneficence”, often over-emphasizing respect for patient autonomy or social justice at the expense of proactively pursuing the patient’s best interests.
Medical domain fine-tuning can significantly improve models’ performance on comprehensive medical ethics tasks, particularly in mitigating weaknesses related to the principle of beneficence, as medical data and expert annotations guide the model to better reflect the intent of beneficence. However, this adaptation risks the model forgetting core ethical knowledge. These findings point toward a future direction: the development of medical LLMs must not only pursue general capabilities but also prioritize ethical alignment to prevent the forgetting of critical ethical principles.
PrinciplismQA provides a robust and scalable framework for diagnosing these deficiencies. Its granular, principle-based analysis can guide targeted improvements, fostering the development of more balanced, context-aware, and responsible AI for healthcare.
Appendix A Data Source

### v2

@@ 3.1 PrinciplismQA Components
Figure 3: Construction workflow of PrinciplismQA.In (a) Data Curation phase, entities highlighted in blue represent GPT-4o. “SOTA LLMs” refers to GPT-4.1, Gemini 2.5 Pro, and Claude 4 Sonnet.
3.2 Data Protocols
Our protocol systematically operationalizes Principlism into structured assessment tasks. We developed a two-stage methodology: First, we mapped principlist concepts to clinical scenarios through a comprehensive taxonomy listed in Table 1, defining 16 ethical dimensions across four principles. Each question in PrinciplismQA is labeled according to these criteria, ensuring philosophical grounding in recognized medical ethics dimensions. Second, we developed a competency-based annotation framework aligned with ACGME Six Core CompetenciesSwing (2007), annotating each rubric item with its corresponding domain, detailed in Appendix D.
We developed a standardized assessment pipeline to evaluate knowledge understanding and principlist reasoning. For Knowledge questions (MCQA), we compare LLM answers against expert-validated ground truth, yielding binary correct/incorrect assessments. For Practice questions (open-ended), we implement rubric-based scoring with multiple expert-defined items per scenario. Figure 2 illustrates this process: given a clinical dilemma, the LLM generates a response, which our Evaluator assesses against predefined rubric items, assigning Partial credit (+0.5) for partially addressed points, Full credit (+1.0) for fully matched reasoning, and No credit (+0.0) for unaddressed considerations. The final score is the sum of gained points divided by total possible points (Gained/Sum), enabling nuanced quantification of ethical reasoning quality.
Data quality was ensured by multi-round curation, calibration, and validation procedures. During these phases, medical experts independently evaluate each LLM response across four dimensions: (1) correctness and preference alignment, (2) clinical relevance, (3) feasibility, and (4) coherence.
3.3 PrinciplismQA Curation, Calibration, and Validation
Figure 3 presents the complete PrinciplismQA construction workflow. For MCQAs, SOTA LLMs classified exam-worthy sections as content of interest (COIs) from 350 international medical ethics textbooks and re-organized the contents to MCQAs. A total of 1,466 medical ethics case analysis articles from the "CASE AND COMMENTARY" section of the AMA Journal of Ethics website were sourced as clinical dilemmas. To ensure the reliability and validity of PrinciplismQA, we implemented a rigorous Human-in-the-Loop verification process involving 12 medical experts (4 practicing physicians and 8 medical postgraduates). LLMs were utilized solely for auxiliary text processing, while all content generation and quality control remained under strict human supervision.
Traceability and Content Fidelity
We ensure full traceability for all dataset components. MCQAs are derived from authoritative textbooks, while open-ended cases originate from the AMA Journal of Ethics. To prevent hallucination during curation, textbook data were first segmented using rule-based matching. GPT-4o was employed strictly to identify concepts suitable for extraction, without transcribing or rewriting the original text. A manual audit of a 10% random sample confirmed a 98.3% accuracy in content preservation.
Rubric Creation and Granularity
For open-ended questions, SOTA LLMs initially extracted candidate keypoints grounded in expert commentaries from the AMA Journal of Ethics. These candidate items were then strictly reviewed, refined, and validated by the expert panel to ensure each keypoint captured a clinically necessary ethical consideration. The number of keypoints per case ranges from 3 to 8, with an average of approximately 4.4 keypoints across the 1,521 open-ended questions. The final score for each question is intrinsically normalized by dividing the gained points by the total possible points for that specific question (Gained/Sum), ensuring that differences in rubric length across cases do not bias aggregate scores.
Difficulty Pre-screening and Quality Filtering
To ensure benchmark difficulty and remove trivially easy, duplicated, or ambiguous items prior to human calibration, we conducted a pre-validation filtering step. Questions that were correctly answered by state-of-the-art models (OpenAI o3 and Gemini 2.5 Flash) were excluded from the dataset. This step served as an initial quality gate, not a substitute for contamination control; conventional training overlap detection was not fully applicable given that several evaluated models (GPT, Claude, Gemini) do not disclose their training corpora. The primary quality gate governing clinical validity remained the model-agnostic human expert calibration process described above. Critically, the resulting dataset does not reflect adversarial selection: post-calibration analysis shows that 558 questions (25.6%) were answered correctly by all evaluated models, while only 47 questions (2.2%) were failed by all, confirming a broad and balanced difficulty distribution.
Inter-Annotator Agreement and Data Quality
For the Knowledge subset, each of the $\sim$2,500 curated MCQAs was independently reviewed by two experts. This process resulted in 393 revisions and 318 deletions, achieving a 96.3% consistency (84.9% in cases of initial disagreement). Ultimately, 87.3% of the MCQAs were retained.
For the Practice subset (1,521 questions, 6,692 keypoints), each item underwent review by at least two experts. Among 274 keypoints where reviewers disagreed, 68.2% were resolved through the physician panel discussion. This protocol yielded a 95.9% overall consistency rate. Notably, 96.4% of the open-ended questions passed the expert review, and only 2.8% of the extracted keypoints required revision.
Evaluation Independence
To prevent circularity between dataset construction and pipeline validation, the three clinical experts who independently graded the 480 question–response pairs in the reliability study (Section 4.4) were completely excluded from the design and annotation of those specific items.
Model Knowledge Practice Overall
General Large Language Model
OpenAI o3 74.4 80.7 77.5
Qwen-Plus 70.0 73.3 71.6
Gemini 2.5 Flash 70.2 72.4 + 71.3 +
OpenAI o3-mini 73.3 67.2 70.2
Claude Sonnet 4
Thinking
70.0 67.5 + 68.7 +
DeepSeek-R1 68.0 + 66.6 + 67.3 +
Gemma3-27B 65.5 40.1 52.8
Gemma3-4B 59.5 42.8 51.1
General Large Reasoning Model
GPT-4.1 74.7 70.8 72.7
Gemini 2.5 Flash
Non-thinking
70.4 + 69.5 69.9
Claude Sonnet 4 70.0 66.6 68.3
Llama-3.1-70B 69.7 55.6 62.6
DeepSeek-V3 66.5 62.5 64.5
Qwen2.5-72B 69.8 53.5 61.7
Qwen2.5-7B 66.4 49.4 57.9
Llama-3.1-8B 58.4 48.5 53.5
Medical LLM/LRM
HuatuoGPT-o1-72B 70.1 61.6 65.9
HuatuoGPT-o1-70B 67.5 61.3 64.4
Med42-70B 67.4 61.2 64.3
MedGemma-27B 64.4 64.3 64.3
HuatuoGPT-o1-7B 66.5 55.2 60.8
HuatuoGPT-o1-8B 55.4 56.4 55.9
MedGemma-4B 59.4 52.9 56.1
Med42-8B 60.5 49.6 55.1
Table 4: Performance of All LLMs on PrinciplismQA. The bold data are the most significant performance in the same category, while the underlined data are the weakest performance. The red color highlights the highest performance and the blue refers to the weakest. “+” denotes stronger performance of reasoning and chat variants within a model family. $\uparrow$” indicates metric improvement of a medical model compared to its general-domain baseline model.
4 Case Studies with PrinciplismQA
@@ 4.1 Experiment Settings
To comprehensively assess Principlism-based ethical reasoning in both general-purpose and domain-specialized language models, we include a broad set of recent medical LLMs alongside general modelsChen et al. (2024); Christophe et al. (2024); Sellergren et al. (2025); Liu et al. (2024). Medical models are fine-tuned for healthcare contexts, allowing us to evaluate whether domain adaptation improves ethical sensitivity and Principlism coverage in clinical scenarios. Besides, representative closed-source LLM families, including ChatGPTOpenAI (2025), Claude 4Anthropic (2025), Qwen3Yang et al. (2025), and Gemini 2.5Comanici et al. (2025), were involved to benchmark the performance of widely used proprietary systems. Commonly used baseline models, LlaMA3.1Dubey et al. (2024), Qwen 2.5Qwen et al. (2025), and GemmaTeam et al. (2025), are evaluated both as general LLMs and as the bases for their corresponding medical model variants, enabling direct comparison between general and medical domain LLMs. All tested LLMs are listed in Table 7 in Appendix B.
All evaluations were conducted using a fixed sampling temperature of 0.1, regardless of whether the model was accessed via API or hosted locally. Each question was tested with a single response per model, with no answer aggregation. Open-source model inference was performed on four NVIDIA H20 GPUs (140GB each), using the original precision as provided by official HuggingFace checkpoints. The prompt constraints and evaluation metrics to be obtained are detailed in Section 3.2.
4.2 Results and Analysis
Model Autonomy Nonmaleficence Beneficence Justice
Claude Sonnet 4
Thinking
0.699 0.681 0.690 0.777 0.675 0.726 0.464 0.682 0.573 0.800 0.659 0.730
Gemini 2.5 Flash
Non-thinking
0.687 0.700 0.694 0.736 0.705 0.720 0.491 0.701 0.596 0.812 0.671 0.741
Table 5: Principlism-Specific Performance of All LLMs on PrinciplismQA.
Takeaway 1: Ethical issues exist for every LLMs.
Takeaway 2: LLMs know ethics but struggles with practice.
Large Reasoning Model vs. Large Language Model in Ethics Across all evaluated models, SOTA closed-source and general reasoning models demonstrated the strongest performance in medical ethics tasks. For example, o3 achieved the highest overall score of 77.5, with 74.4% Knowledge accuracy and an 80.7 performance on Practice, while GPT-4.1 led among chat models with an overall score of 72.7, both outperforming cutting-edge Practice performance and specialized medical models. Reasoning-focused variants, such as gemini-2.5-flash, claude-sonnet-4, and deepseek, consistently surpassed their chat-oriented counterparts in Practice scenarios. These results suggest that models with stronger foundational and reasoning capabilities are better equipped to handle complex, non-standardized ethical dilemmas in the medical domain.
Takeaway 3: Reasoning helps ethics.
4.3 Fine-grained Analysis
Figure 4: Competency-specific open-ended question performance comparisons: (a) by model category, (b) between medical LLMs and their baseline models. “ICS.”, “MK.”, “PBLI.”, “PC.”, “Pf”, and “SBP.” are the abbreviations of “Interpersonal and Communication Skills”, “Medical and Ethical Knowledge”, “Practice-Based Learning and Improvement”, “Patient Care”, “Professionalism”, and “Systems-Based Practice”.
Takeaway 5: LLMs struggle most with beneficence; fine-tuning helps.
4.4 Trustworthiness of the Assessment
To validate the effectiveness of our automated assessment pipeline, we conducted a reliability study comparing its scoring against human expert consensus. We sampled 480 question–response pairs (covering 1,516 individual keypoints) and employed three clinical experts to grade them independently.
As shown in Table 6, the inter-rater reliability (ICC) among the three human experts was 0.67, reflecting the inherent subjectivity and difficulty in grading open-ended ethical reasoning. In comparison, the ICC between our Assessment Pipeline and the mean score of the human experts reached 0.71. This result indicates that our pipeline not only achieves grading consistency comparable to human experts but slightly surpasses the average human consensus. This validates that our expert-calibrated pipeline serves as a scalable, consistent, and highly reliable evaluator for complex medical ethics assessments. We also provide score scatter plots comparing human experts and our Assessment Pipeline in Appendix E.
Grader Comparison ICC
Human–Human (3 experts) 0.67
Assessment Pipeline vs. Human Mean 0.71
Table 6: Inter-rater reliability (ICC) comparison between Human Experts and our Assessment Pipeline.
5 Conclusion
We introduced PrinciplismQA, a philosophy-grounded approach addressing the critical gap between medical LLMs’ knowledge accuracy and ethical reasoning capability in clinical contexts. By systematically incorporating Principlism into assessment design through expert-validated protocols, our approach enables reproducible evaluation of LLM ethical alignment at scale. Our case studies on recent LLMs demonstrate that high performance on knowledge benchmarks does not translate to ethical considerations in clinical decision-making, revealing substantial gaps in navigating ethical trade-offs across multiple valid solutions. PrinciplismQA provides the research community and industry with a validated methodology for assessing clinical AI deployment readiness, bridging the gap between technological capability and ethics trustworthiness. Future work should extend this approach to multi-modal clinical scenarios and investigate methods for improving LLM principlist reasoning through targeted training interventions.
Limitations
PrinciplismQA opens several directions for future work. First, our benchmark is currently text-only, while real-world clinical decisions often involve multimodal information such as medical images, patient charts, and vital signs. Extending to multimodal scenarios would enable more comprehensive ethical reasoning assessment in realistic clinical contexts. Second, with 3,648 questions, PrinciplismQA is designed for evaluation rather than training. Scaling the dataset through our expert-calibrated protocol could support targeted fine-tuning to improve LLM principlist reasoning capabilities, for instance through training on principlist reasoning tasks. Third, our assessment pipeline relies on an LLM-as-a-Judge scoring module, which is practically necessary for evaluating open-ended ethical reasoning at scale. Our reliability study demonstrates that the pipeline achieves ICC of 0.71 against human expert consensus, surpassing inter-human agreement of 0.67, yet future evaluators could explore hybrid approaches that further integrate expert judgment to address edge cases and potential conflation of response fluency with reasoning quality. Fourth, while Principlism serves as an internationally recognized gold standard in clinical ethics, future work should examine cross-cultural variations in ethical norms and incorporate diverse regional case sources to broaden the benchmark’s global applicability. Finally, PrinciplismQA targets LLMs in decision-support roles that surface ethical considerations for clinicians rather than autonomous ethical decision-making. Future work should investigate whether principlist reasoning performance correlates with improved human-AI collaborative outcomes in realistic deployment workflows such as scribing, documentation, and care coordination.
Acknowledgments
This work is supported by Longgang District Special Funds for Science and Technology Innovationunder Grant LGKCSDPT2025002.
Appendix A Data Scope
The scope of PrinciplismQA encompasses high-quality multiple-choice questions derived from authoritative medical ethics textbooks. To ensure content fidelity, each item preserves the original source context. Figure 5 illustrates a representative sample of a curated MCQA, displaying the mapping from the Content of Interest (COI) to the structured question, options, and expert-verified explanation.
Appendix B Candidate LLMs
To ensure a comprehensive evaluation of ethical reasoning across different model architectures and training paradigms, we selected a diverse set of candidate models ranging from general-purpose LLMs to specialized medical models. This selection enables a direct comparison between general reasoning capabilities and domain-specific adaptation in the context of clinical ethics. Table 7 provides the complete list of evaluated models along with their access details.
Model API Provider / HF Checkpoint
@@ Appendix B Candidate LLMs
Appendix C Data Source